News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 26 August 2005 11.00am EDT

During the last week the observing schedule was interrupted following a high radiation event at 3:40pm EDT on Aug 22 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and radiation levels dropped sufficiently to for the loads to be restarted at 6.38pm EDT on Aug 24. However, SCS 107 ran again shortly after loads were restarted during a LETG insertion that resulted in the grating remaining partially inserted. The second execution of SCS 107 was determined not to be due to high radiation and is under investigation. Preliminary findings suggest the radmon process triggered on stale data due to a specific timing condition.

A real-time procedure was completed on Aug 25 to retract the LETG in preparation for resuming science. In addition, procedures were completed to re-enable the sun-position and momentum monitors, and to perform a maneuver to ensure a favorable thermal attitude.

A realtime procedure was conducted on Aug 19 to uplink and verify a corrected patch for the EPHIN Off capability. In the event of an EPHIN failure due to elevated thermal conditions, the patch will allow the HRC event and anti-coincidence shield rates to be used by the radiation monitor process to monitor high radiation events.

------------------------------------------
The schedule of targets for the next week has yet to be finalized.

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports